OpenStudy (anonymous):

The slope of a line that is perpendicular to a line whose slope is -2/7 is

OpenStudy (ccswims):

since it's perpendicular to switch it

OpenStudy (ccswims):

7/2

TheSmartOne (thesmartone):

Perpendicular slopes when multiplied equal to -1 So lets say the slope of the perpendicular line is x

TheSmartOne (thesmartone):

\[\frac{ -2 }{ 7 }\times x=-1\]

TheSmartOne (thesmartone):

\[\frac{ -7 }{ 2 } \times \frac{ -2 }{ 7 }\times x=-1 \times \frac{ -7 }{ 2 }\]

TheSmartOne (thesmartone):

\[x=\frac{ 7 }{ 2 }\]

TheSmartOne (thesmartone):

@tommtwo